# Solitary cluster waves in periodic potentials: Formation, propagation, and soliton-mediated particle transport

### v1

@@ Abstract
Transport processes in crowded periodic structures are often mediated by cooperative movements of particles forming clusters. Recent theoretical and experimental studies of driven Brownian motion of hard spheres showed that cluster-mediated transport in one-dimensional periodic potentials can proceed in form of solitary waves. We here give a comprehensive description of these solitons. Fundamental for our analysis is a static presoliton state, which is formed by a periodic arrangements of basic stable clusters. Their size follows from a geometric principle of minimum free space. Adding one particle to the presoliton state gives rise to solitons. We derive the minimal number of particles needed for soliton formation, number of solitons at larger particle numbers, soliton velocities and soliton-mediated particle currents. Incomplete relaxations of the basic clusters are responsible for an effective repulsive soliton-soliton interaction seen in measurements. Our results provide a theoretical basis for describing experiments on cluster-mediated particle transport in periodic potentials.
I Introduction
Particle transport in densely populated periodic structures frequently proceeds by cooperative movements of particle assemblies. Examples are surface crowdions in copper adatom diffusion on a strained surface [1], interstitial crowdions in crystals [2, 3], and similar types of vacancy clusterings in form of anti-crowdions [4] or voidions [5, 6]. A further example are defect motions in colloidal monolayers, which are completely filled wells of a two-dimensional periodic or quasi-periodic optical potential [7, 8, 9, 10, 11, 12]. Under time-dependent forcing, cluster-mediated transport was seen for vortices in a nanostructured superconducting film driven by an oscillating electric current [13], paramagnetic colloids above a magnetic bubble lattice driven by a rotating magnetic field [14, 15, 16], and polystyrene particles driven by an oscillating periodic optical potential [17].
@@ I Introduction
Figure 1: Illustration of mechanically stable presoliton state, running state carrying solitons, dynamical phase transition between presoliton and running state, and behavior of soliton-mediated particle currents. In (a), the presoliton state in the tilted periodic potential $U(x)-fx$ is depicted. It is formed by a periodic sequence of basic mechanically stable clusters of size $n_{\rm b}=3$ and one residual particle ($n_{\rm res}=1$). In (b), a particle configuration in the running state is displayed, which forms after adding one particle to the presoliton state. One soliton cluster with five particles can be identified (spheres marked in green). The remaining (spheres marked in red) form basic $n_{\rm b}$-clusters. The configurations in (a) and (b) represent a small section of a system with $L=75$ potential wells. How the numbers $N_{\rm sol}$ and $N_{\rm b}$ of solitons and basic stable clusters vary with the overfilling $\Delta$ is shown in (c). $N_{\rm sol}$ jumps from zero to one in the transition between presoliton and running state, and increases linearly with slope two upon further increase of $\Delta$ (two further solitons per added particle). Part (d) shows the particle current mediated by the solitons as a function of $N_{\rm sol}$. For large $N_{\rm sol}$, it is not proportional to $N_{\rm sol}$ due to an effective repulsive interaction between solitons. This effective interaction is reflecting a slowing down of mean soliton velocities with increasing $N_{\rm sol}$, which is shown in the inset of (d). Results in (a)-(d) are obtained from simulations based on Eq. (6) with $U(x)$ from Eq. (4), hard sphere diameter $\sigma=0.6$, and drag force $f=0.1f_{\rm c}$ [$U_{0}=1$, $\lambda=1$, $\mu=1$, and $f_{\rm c}$ is the critical force for overtilting, see Eq. (5)]. They agree with analytical results derived in Secs. IV-X.
@@
II Solitary cluster waves
We consider particles dragged by a constant force $f$ across a sinusoidal potential, for which first experimental observations of the solitary waves were reported in Ref. [21]. The particles’ motion is described by the Langevin equations
@@ II Solitary cluster waves
The dynamics of the particles is constrained to a finite interval of size $L$ with periodic boundary conditions, where $L$ is an integer multiple of $\lambda$. We set $\mu=1$ in the following, and use $U_{0}$, $\lambda$ and $\lambda^{2}/\mu U_{0}$ as units of energy, length and time, respectively.
In our previous work [20], we treated a special case of particle diameters in the range $10.4\leq\sigma<$ 1 for overfilling $\Delta=N-L=1$, i.e. when the number $N$ of particles exceeds the number $L$ of potential wells by one. For high barriers $U_{0}\gg k_{\rm\scriptscriptstyle B}T$ of the periodic potential and weak drag force $f\ll f_{\rm c}$, we found that particle currents as a function of $\sigma$ exhibit peaks at magic particle diameters $\sigma_{n}=(n-1)/n$, $n=2,3,\ldots$ Solitary cluster waves, which manifest themselves as periodic sequences of cluster movements, are responsible for this striking behavior. Their properties could be derived by considering the limit of zero noise, where Eqs. (2) are
@@
We first give in Sec. III basic features of particle clusters in a tilted periodic potential. In Sec. IV we discuss presoliton states, which are mechanically stable states with largest number of particles. The presoliton state is formed by basic mechanically stable clusters composed of $n_{\rm b}$ particles. The knowledge of $n_{\rm b}$ allows us to derive in Sec. V the minimal overfilling necessary to generate solitons. In Sec. VI we describe how solitons propagate by periodic sequences of cluster movements and derive their time period and velocity. Thereafter we analyze in Sec. VII how many solitons form for a given overfilling and discuss in Sec. VIII how our results can be tested in experiments. We explain the effective repulsive soliton-soliton interaction in Sec. IX and derive the particle currents mediated by solitons in Sec. X.
Our analysis is performed for hard-sphere diameters $10<\sigma<$ 1. Section XI discusses how results for $\sigma>1$ larger than the wavelength $\lambda=1$ are obtained from those for $10<\sigma<$ 1.
@@
III Particle clusters in tilted periodic potential
@@
If they are satisfied in an interval $[y,y^{\prime}]$, the time for the $n$-cluster to move from $y$ to $y^{\prime}$ is
@@ IV Presoliton states
In Sec. IV.1, we first derive $n_{\rm b}$ for the case of infinitesimal force $f=0^{+}$, which we refer to as zero-force limit. In this limit, solitons can occur only if the particle diameter $\sigma$ is a rational number (in units of the wavelength $\lambda=1)$,
@@
This is a necessary condition, since $U_{q}(x)=0$ for such $\sigma$, see Eq. (7), i.e. a $q$-cluster could move without surmounting barriers. To represent each such $\sigma_{p,q}$ uniquely, $p$ and $q$ are taken to be coprime. Because $\sigma_{p,q}<1$, it is $q>p$.
The necessary condition $\sigma_{p,q}=p/q$ does not imply that solitons must occur for all $p/q$. For example, in Ref. [20] solitons were studied for overfilling $\Delta=1$ and appeared for $p=q-1$ only. We will show that solitons occur for most $p/q$, if the overfilling exceeds a minimal value $\Delta_{\rm min}$. This $\Delta_{\rm min}$ The required minimal overfilling is calculated in Sec. V. It may, however, not be realizable for a finite system size $L$.
The knowledge for $f=0^{+}$ allows us to determine $n_{\rm b}$ for finite $f<f_{\rm c}$ also, which we show in Sec. IV.2.
IV.1 Basic stable cluster size $n_{\rm b}$ for $f=0^{+}$
@@
The relation between stabilizability and residual free space is given by the following free-space theorem, derived in Appendix A: If an isolated $n$-cluster is stabilizable, its residual free space $r_{n}$ is smaller than that of any cluster of smaller size, i.e. it holds
@@
$n_{\rm b}(\sigma_{p,q})=\operatorname*{argmin}_{n\in\{1,\ldots,q-1\}}(\lceil n\sigma_{p,q}\rceil-n\sigma_{p,q})\,.$ (19)
Equivalently, $n_{\rm b}$ is determined by Eq. (18), i.e. $r_{n_{\rm b}}=\lceil n_{\rm b}\sigma_{p,q}\rceil-n_{\rm b}\sigma_{p,q}=1/q$. This condition can be rewritten in the form
@@
IV.2 Basic stable cluster size $n_{\rm b}$ for $f>0$
@@ V Minimal overfilling
$N_{\rm pre}^{\rm b}=\left\lfloor\frac{L}{\lceil n_{\rm b}\sigma\rceil}\right\rfloor\,.$ (27)
The number of potential wells accommodating all $n_{\rm b}$-clusters is
@@
VI Soliton propagation
VI.1 Propagation modes
@@
In what follows, we focus on the basic propagation mode A. Quantities like soliton potentials or velocities discussed below can be treated analogously for the variant B. For the velocity of type B solitons, we give the corresponding calculation in Appendix B.
@@
Figure 6: (a) Tilted potentials $U_{n_{\rm c}+n_{\rm b}}(x)-fx$ (green) and $U_{n_{\rm c}}-fx$ (orange) of the composite $(n_{\rm c}+n_{\rm b})$- and core $n_{\rm c}$-cluster for $\sigma=0.57$ and $f=0.01$, where $n_{\rm b}=5$ and $n_{\rm c}=2$. The inset shows the soliton potential (36) in the interval $[y_{1}+n_{\rm b}\sigma,y_{1}+\lceil n_{\rm b}\sigma\rceil[$, which is marked by the vertical dashed lines in the main figure. (b) Tilted soliton potential and corresponding force field in the continuous-coordinate representation (see text). For the soliton potential and force fields, the parts stemming from the composite and core soliton cluster are marked in green and orange, respectively.
VI.2 Soliton potential and force field
@@
Taking the soliton position as that of the soliton clusters, i.e. the position $x$ of the leftmost particle of the $n_{\rm c}$- and $(n_{\rm c}+n_{\rm b})$-clusters in the basic propagation mode, we can define a potential for the soliton motion for $x\in[y_{1}+n_{\rm b}\sigma,y_{1}+\lceil n_{\rm b}\sigma\rceil[$: {linenomath*}
$U_{\rm sol}(x)=$ $\left[U_{n_{\rm c}}(x)-U_{n_{\rm c}}(y_{2})\right]\Theta(y_{2}-x)$
${}+\left[U_{n_{\rm c}+n_{\rm b}}(x)-U_{n_{\rm c}+n_{\rm b}}(y_{2})\right]\Theta(x-y_{2})\,.$ (36)
@@
Here, $\Theta(.)$ is the Heaviside step function [$\Theta(y)=1$ for $y\geq 1$ and zero otherwise]. The constants $U_{n_{\rm c}}(y_{2})$ and $U_{n_{\rm c}+n_{\rm b}}(y_{2})$ were subtracted from the cluster potentials to make the soliton potential continuous at $y=y_{2}$. That way, the force $f-U^{\prime}_{\rm sol}(x)$ on the soliton jumps at $y=y_{2}$ from $f-U^{\prime}_{n_{\rm c}}(y_{2})$ to $f-U^{\prime}_{n_{\rm c}+n_{\rm b}}(y_{2})$, but has no $\delta$-function singularity. The tilted soliton potential $U_{\rm sol}(x)-fx$ is shown in the inset of the figure, where the parts stemming from the core and composite soliton clusters are marked as in the main figure.
@@
When a soliton starting at $y_{1}+n_{\rm b}\sigma$ has moved one time period $\tau_{\rm sol}$, its position jumps from $(y_{1}+\lceil n_{\rm b}\sigma\rceil)$ to $y_{1}+\lceil n_{\rm b}\sigma\rceil+n_{\rm b}\sigma$. After the jump, the potential is the same as at the starting position. For displaying the soliton potential, it is helpful to use a continuous-coordinate representation also with a coordinate $x^{\prime}$, where the jumps in the real soliton coordinate are removed. In this representation, the soliton potential is periodic with a period length equal to the residual free space $r_{n_{\rm b}}=\lceil n_{\rm b}\sigma\rceil-n_{\rm b}\sigma$ of the basic stable cluster.
@@
Figure 6(b) shows the tilted soliton potential $U_{\rm sol}(x)-fx$ from the inset of Fig. 6(a) in the continuous-coordinate representation, and the corresponding force field $f-U^{\prime}_{\rm sol}(x)$. The force is always positive and it jumps when a core cluster attaches to a composite cluster. This corresponds to changes of the color from orange to green. When the color changes from green to orange, a core cluster detaches from a composite cluster. In such detachments events, the real soliton coordinate jumps by $n_{\rm b}\sigma$. These jumps are not visible in the continuous-coordinate representation in Fig. 6(b).
VI.3 Soliton velocity

### v2

@@ Abstract
Transport processes in crowded periodic structures are often mediated by cooperative movements of particles forming clusters. Recent theoretical and experimental studies of driven Brownian motion of hard spheres showed that cluster-mediated transport in one-dimensional periodic potentials can proceed in form of solitary waves. We here give a comprehensive description of these solitons. Fundamental for our analysis is a static presoliton state, which is formed by a periodic arrangement of basic stable clusters. Their size follows from a geometric principle of minimum free space. Adding one particle to the presoliton state gives rise to solitons. We derive the minimal number of particles needed for soliton formation, number of solitons at larger particle numbers, soliton velocities and soliton-mediated particle currents. Incomplete relaxations of the basic clusters are responsible for an effective repulsive soliton-soliton interaction seen in measurements. A dynamical phase transition is predicted to occur in current-density relations at low temperatures. Our results provide a theoretical basis for describing experiments on cluster-mediated particle transport in periodic potentials.
journal: Chaos, Solitons & Fractals
[…]
[inst1]organization=Universität Osnabrück, Institut für Physik, Fachbereich Mathematik/Informatik/Physik, addressline=Barbarastraße 7, postcode=D-49076, city=Osnabrück, country=Germany
[…]
[inst2]organization=Institut für Theoretische Physik II: Weiche Materie, Heinrich-Heine-Universität Düsseldorf,
addressline=Universitätsstraße 1,, postcode=D-40225, city=Düsseldorf, country=Germany
[…]
[inst3]organization=Charles University, Faculty of Mathematics and Physics, Department of Macromolecular Physics, addressline=V Holešovičkách 2, postcode=
CZ-18000, city=Prague, country=Czech Republic
1 Introduction
Particle transport in densely populated periodic structures frequently proceeds by cooperative movements of particle assemblies. Examples are surface crowdions in copper adatom diffusion on a strained surface [1], and interstitial crowdions in crystals [2, 3, 4]. These crowdions are low-temperature configurations of interstitial atoms densely packed in one direction, which can extend over many lattice constants. Analogous clusterings of vacancies can occur, which are referred to as anti-crowdions [5] or voidions [6, 7]. A further example are defect motions in colloidal monolayers, which are completely filled wells of a two-dimensional periodic or quasi-periodic optical potential [8, 9, 10, 11, 12, 13]. Under time-dependent forcing, cluster-mediated transport was seen for vortices in a nanostructured superconducting film driven by an oscillating electric current [14], paramagnetic colloids above a magnetic bubble lattice driven by a rotating magnetic field [15, 16, 17], and polystyrene particles driven by an oscillating periodic optical potential [18].
@@ 1 Introduction
Figure 1: Illustration of mechanically stable presoliton state, running state carrying solitons, dynamical phase transition between presoliton and running state, and behavior of soliton-mediated particle currents. In (a), the presoliton state in the tilted periodic potential $U(x)-fx$ is depicted. It is formed by a periodic sequence of basic mechanically stable clusters of size $n_{\rm b}=3$ and one residual particle ($n_{\rm res}=1$). In (b), a particle configuration in the running state is displayed, which forms after adding one particle to the presoliton state. One soliton cluster with five particles can be identified (spheres marked in green). The remaining (spheres marked in red) form basic $n_{\rm b}$-clusters. The configurations in (a) and (b) represent a small section of a system with $L=75$ potential wells. How the numbers $N_{\rm sol}$ and $N_{\rm b}$ of solitons and basic stable clusters vary with the overfilling $\Delta$ is shown in (c). $N_{\rm sol}$ jumps from zero to one in the transition between presoliton and running state, and increases linearly with slope two upon further increase of $\Delta$ (two further solitons per added particle). Part (d) shows the particle current mediated by the solitons as a function of $N_{\rm sol}$. For large $N_{\rm sol}$, it is not proportional to $N_{\rm sol}$ due to an effective repulsive interaction between solitons. This effective interaction is reflecting a slowing down of mean soliton velocities with increasing $N_{\rm sol}$, which is shown in the inset of (d). Results in (a)-(d) are obtained from simulations based on Eq. (6) with $U(x)$ from Eq. (4), hard sphere diameter $\sigma=0.6$, and drag force $f=0.1f_{\rm c}$ [$U_{0}=1$, $\lambda=1$, $\mu=1$, and $f_{\rm c}$ is the critical force for overtilting, see Eq. (5)]. They agree with analytical results derived in Secs. 4-10.
@@
2 Solitary cluster waves
In general, solitons are waves that travel without dispersion and maintain their shape [24, 20, 25]. They were modeled and observed across various applications, including matter waves in Bose-Einstein condensates [26, 27], optical waves in Kerr frequency combs [28, 29], Alfvén magnetic waves in plasmas [30, 31], compression waves in damped Newtonian dynamics of Toda lattice chains subject to time-periodic driving [32, 33, 34, 35], waves of stable defects in periodically inhomogeneous long Josephson junctions called supersolitons [36] or superfluxons [37], as well as waves in large-scale phenomena such as tsunamis [38, 39] and tidal bores [40].
Here we consider particles dragged by a constant force $f$ across a sinusoidal potential, for which first experimental observations of the solitary waves were reported in Ref. [22].
The particles perform an overdamped Brownian motion, where effects of inertia are negligible. This motion is described by the Langevin equations [41]
@@ 2 Solitary cluster waves
The dynamics of the particles is constrained to a finite interval of size $L$ with periodic boundary conditions, where $L$ is an integer multiple of $\lambda$.
In our previous work [21], we treated a special case of particle diameters in the range $0.4\,\lambda\leq\sigma<\lambda$ for overfilling $\Delta=N-L=1$, i.e. when the number $N$ of particles exceeds the number $L$ of potential wells by one. For high barriers $U_{0}\gg k_{\rm\scriptscriptstyle B}T$ of the periodic potential and weak drag force $f\ll f_{\rm c}$, we found that particle currents as a function of $\sigma$ exhibit peaks at magic particle diameters $\sigma_{n}/\lambda=(n-1)/n$, $n=2,3,\ldots$ Solitary cluster waves, which manifest themselves as periodic sequences of cluster movements, are responsible for this striking behavior. Their properties could be derived by considering the limit of zero noise, where Eqs. (2) are
@@
In Sec. 3 we first give basic features of particle clusters in a tilted periodic potential. In Sec. 4 we discuss presoliton states, which are mechanically stable states with largest number of particles. The presoliton state is formed by basic mechanically stable clusters composed of $n_{\rm b}$ particles. The knowledge of $n_{\rm b}$ allows us to derive in Sec. 5 the minimal overfilling necessary to generate solitons. In Sec. 6 we describe how solitons propagate by periodic sequences of cluster movements and derive their time period and velocity. Thereafter we analyze in Sec. 7 how many solitons form for a given overfilling and discuss in Sec. 8 how our results can be tested in experiments. We explain the effective repulsive soliton-soliton interaction in Sec. 9 and derive the particle currents mediated by solitons in Sec. 10.
We use $U_{0}$, $\lambda$ and $\lambda^{2}/\mu U_{0}$ as units of energy, length and time in the following. Our analysis is performed for hard-sphere diameters $10<\sigma<$ 1. Section 11 discusses how results for $\sigma>1$ larger than the wavelength $\lambda=1$ are obtained from those for $10<\sigma<$ 1.
@@
3 Particle clusters in tilted periodic potential
@@
If they are satisfied in an interval $[y,y^{\prime}]$ of positions, the time for the $n$-cluster to move from $y$ to $y^{\prime}$ is
@@ 4 Presoliton states
In Sec. 4.1, we first derive $n_{\rm b}$ for the case of infinitesimal force $f=0^{+}$, which we refer to as zero-force limit. In this limit, solitons can occur only if the particle diameter $\sigma$ is a rational number (in units of the wavelength $\lambda=1)$,
@@
This is a necessary condition, since the cluster potential $U_{n}(x)$ becomes zero if $n=q$ and $\sigma=p/q$. It means that a $q$-cluster composed of particles of size $\sigma_{p,q}$ can move without surmounting barriers. To represent each such $\sigma_{p,q}$ uniquely, $p$ and $q$ are taken to be coprime. Because $\sigma_{p,q}<1$, it is $q>p$.
The necessary condition $\sigma_{p,q}=p/q$ does not imply that solitons must occur for all $p/q$. For example, in Ref. [21] solitons were studied for overfilling $\Delta=1$ and appeared for $p=q-1$ only. We will show that solitons occur for most $p/q$, if the overfilling exceeds a minimal value $\Delta_{\rm min}$. This $\Delta_{\rm min}$ The required minimal overfilling is calculated in Sec. 5. It may, however, not be realizable for a finite system size $L$.
The knowledge for $f=0^{+}$ allows us to determine $n_{\rm b}$ for finite $f<f_{\rm c}$ also, which we show in Sec. 4.2.
4.1 Basic stable cluster size $n_{\rm b}$ for $f=0^{+}$
@@
The relation between stabilizability and residual free space is given by the following free-space theorem, derived in A: If an isolated $n$-cluster is stabilizable, its residual free space $r_{n}$ is smaller than that of any cluster of smaller size, i.e. it holds
@@
$n_{\rm b}(\sigma_{p,q})=\operatorname*{argmin}_{n\in\{1,\ldots,q-1\}}(\lceil n\sigma_{p,q}\rceil-n\sigma_{p,q})\,,$ (19)
where $\operatorname*{argmin}(f(x))$ gives the argument $x$ at which the function $f(x)$ has its minimum. Hence, $n_{\rm b}(\sigma_{p,q})$ is equal to that $n\in\{1,\ldots,q-1\}$, where $f(n)=\lceil n\sigma_{p,q}\rceil-n\sigma_{p,q}$ is minimal. Equivalently, $n_{\rm b}$ is determined by Eq. (18), i.e. $r_{n_{\rm b}}=\lceil n_{\rm b}\sigma_{p,q}\rceil-n_{\rm b}\sigma_{p,q}=1/q$. This condition can be rewritten in the form
@@
4.2 Basic stable cluster size $n_{\rm b}$ for $f>0$
@@ 5 Minimal overfilling
$N_{\rm pre}^{\rm b}=\left\lfloor\frac{L}{\lceil n_{\rm b}\sigma\rceil}\right\rfloor\,,$ (27)
where $\lfloor x\rfloor$ is the largest integer smaller than $x$. The number of potential wells accommodating all $n_{\rm b}$-clusters is
@@
6 Soliton propagation
6.1 Propagation modes
@@
In what follows, we focus on the basic propagation mode A. Quantities like soliton potentials or velocities discussed below can be treated analogously for the variant B. For the velocity of type B solitons, we give the corresponding calculation in B.
@@
6.2 Soliton potential and force field
@@
Taking the soliton position as that of the soliton clusters, i.e. the position $x$ of the leftmost particle of the $n_{\rm c}$- and $(n_{\rm c}+n_{\rm b})$-clusters in the basic propagation mode, we can define a potential for the soliton motion for $y\in[y_{1}+n_{\rm b}\sigma,y_{1}+\lceil n_{\rm b}\sigma\rceil[$:
$U_{\rm sol}(y)=$ $\left[U_{n_{\rm c}}(y)-U_{n_{\rm c}}(y_{2})\right]\Theta(y_{2}-y)$
${}+\left[U_{n_{\rm c}+n_{\rm b}}(y)-U_{n_{\rm c}+n_{\rm b}}(y_{2})\right]\Theta(y-y_{2})\,.$ (36)
@@
Here, $\Theta(.)$ is the Heaviside step function [$\Theta(y)=1$ for $y\geq 1$ and zero otherwise]. The constants $U_{n_{\rm c}}(y_{2})$ and $U_{n_{\rm c}+n_{\rm b}}(y_{2})$ were subtracted from the cluster potentials to make the soliton potential continuous at $y=y_{2}$. That way, the force $f-U^{\prime}_{\rm sol}(y)$ on the soliton jumps at $y=y_{2}$ from $f-U^{\prime}_{n_{\rm c}}(y_{2})$ to $f-U^{\prime}_{n_{\rm c}+n_{\rm b}}(y_{2})$, but has no $\delta$-function singularity. The tilted soliton potential $U_{\rm sol}(y)-fy$ is shown in the inset of Fig. 6(a), where the parts stemming from the core and composite soliton clusters are marked as in the main figure.
When a soliton starting at $y_{1}+n_{\rm b}\sigma$ has moved one time period $\tau_{\rm sol}$, its position jumps from $(y_{1}+\lceil n_{\rm b}\sigma\rceil)$ to $y_{1}+\lceil n_{\rm b}\sigma\rceil+n_{\rm b}\sigma$, because when the $(n_{\rm c}+n_{\rm b})$-cluster reaches position $(y_{1}+\lceil n_{\rm b}\sigma\rceil)$, an $n_{c}$-cluster at its right end detaches, which then becomes the soliton cluster at position $y_{1}+\lceil n_{\rm b}\sigma\rceil+n_{\rm b}\sigma$, see Fig. 6(a). After the jump of size $n_{\rm b}\sigma$, the potential $U_{\rm sol}$ is the same as at the starting position.
Figure 6: (a) Tilted potentials $U_{n_{\rm c}+n_{\rm b}}(x)-fx$ (green) and $U_{n_{\rm c}}-fx$ (orange) of the composite $(n_{\rm c}+n_{\rm b})$- and core $n_{\rm c}$-cluster for $\sigma=0.57$ and $f=0.01$, where $n_{\rm b}=5$ and $n_{\rm c}=2$. The clusters of sizes $n_{\rm b}$ and $(n_{\rm c}+n_{\rm b})$ have positions in the intervals bounded by the vertical dashed lines. When an $(n_{\rm c}+n_{\rm b})$-cluster reaches position $y_{1}+\lceil n_{\rm b}\sigma\rceil$, an $n_{\rm c}$-cluster detaches at its right end and the soliton propagation continues with this $n_{\rm c}$-cluster at position $y_{1}+\lceil n_{\rm b}\sigma\rceil+n_{\rm b}\sigma$. The corresponding jump in soliton position of size $n_{\rm b}\sigma$ is indicated by the horizontal arrow. The inset shows the soliton potential (36) in the interval $[y_{1}+n_{\rm b}\sigma,y_{1}+\lceil n_{\rm b}\sigma\rceil[$. (b) Tilted soliton potential $U_{\rm sol}(y)-fy$ and corresponding force field in the continuous-coordinate representation. For this representation, the potential $U_{\rm sol}(y)$ from the interval $[y_{1}+n_{\rm b}\sigma,y_{1}+\lceil n_{\rm b}\sigma\rceil[$ in (a) is periodically continued. Parts stemming from the composite and core soliton cluster are marked in green and orange, respectively.
@@
For displaying the soliton potential, it is helpful to use a continuous-coordinate representation where the jumps in the real soliton coordinate are removed. This means that we periodically continue $U_{\rm sol}(y)$ from the interval $y\in[y_{1}+n_{\rm b}\sigma,y_{1}+\lceil n_{\rm b}\sigma\rceil[$ in Eq. (36). The soliton potential in the continuous coordinate representation hence is periodic with a period length equal to the residual free space $r_{n_{\rm b}}=\lceil n_{\rm b}\sigma\rceil-n_{\rm b}\sigma$ of the basic stable cluster.
@@
Figure 6(b) shows the tilted soliton potential $U_{\rm sol}(y)-fy$ for three periods $r_{n_{\rm b}}$ of $U_{\rm sol}(y)$, in the continuous-coordinate representation, and the corresponding force field $f-U^{\prime}_{\rm sol}(y)$. The force is always positive and it jumps when a core cluster attaches to a composite cluster. This corresponds to changes of the color from orange to green. When the color changes from green to orange, a core cluster detaches from a composite cluster. In such detachments events, the real soliton coordinate jumps by $n_{\rm b}\sigma$. These jumps are not visible in the continuous-coordinate representation in Fig. 6(b).
6.3 Soliton velocity